The hydrogen sulfide in a 50.0-g sample of crude p
Answer #1

In this process the sulfur (S) in the form of H2S present in the sample is finally trapped in 0.108 g of CdSO4

Step1:

Calculate the # moles of CdSO4 = mass of CdSO4/mol wt of CdSO4 = 0.108/208 = 0.00053 moles

these are also the # moles of S which are the same as the # moles of H2S

Step 2:

From the moles of H2S find the mass

mass of H2S = moles * mol wt H2S = 0.00053 * 34 = 0.018 g

Step 3:

Weight of the original sample is 50.0 g

% H2S = 0.018/50.0 * 100 = 0.036%